# Improving Technical ”How-to” Query Accuracy with Automated Search Results Verification and Reranking

### v1

@@ Abstract.
Many people use search engines to find online guidance to solve computer or mobile device problems. Users frequently encounter challenges in identifying effective solutions from search results, often wasting time trying ineffective solutions that seem relevant yet fail to solve the real problems. This paper introduces a novel approach to improving the accuracy and relevance of online technical support search results through automated search results verification and reranking. Taking ”How-to” queries specific to on-device execution as a starting point, we first developed a solution that allows an AI agent to interpret and execute step-by-step instructions in the search results in a controlled Android environment. We further integrated the agent’s findings into a reranking mechanism that orders search results based on the success indicators of the tested solutions.
@@ 1. Introduction
Often people lack the knowledge to complete a task and rely on a search engine to retrieve potential instructions by asking ”How-to” or similar style queries. These requests are ”procedural queries” where the user’s intent is to retrieve step-by-step instructions for performing a specific task or operation. However, instructions extracted from retrieved results may not be executable, as the search engine primarily ranks pages based on similiarty of search query and candidate web page (Qin and Liu, 2013). Furthermore, although user engagement information, such as search logs and page quality indicators like PageRank, are utilized for ranking results in modern search engines, the optimal solution often varies on a user-specific basis. The use of different operating systems, screen types, and app versions motivates a personalized reranker. In addition, search engines integrated with large language models (LLMs) haven shown to involve the likelihood of producing hallucinations (Lewis et al., 2020).
@@
How do users solve the problem? Typically, a user could visit multiple retrieved pages. If a page that looks relevant contains some step-by-step instructions, the user might try to follow the instructions to see whether it solves the problem. However, users frequently experience frustration and waste considerable time attempting various instructions across different pages. Can a search engine assist the user in a more efficient way? Motivated by the users’ time-consuming manual solution, we substitute this tedious process with an automatic instruction verification process as an additional component of the search process. Our hypothesis is that search results verification for technical ”How-to” queries can achieve decent accuracy given the recent research progress on multimodal LLMs, especially with the help of state-of-the-art GPT agents.
@@
In particular, we introduce a three-stage solution to do instruction set verification and reranking relevant pages, as illustrated in Figure 1: The first stage involves an Instruction Extraction Model to obtain step-by-step instructions from each retrieved page for a ”How-to” search query. The second stage verifies their quality by automatically trying or simulating the instruction execution on devices. The third stage reranks retrieved results based on the execution information.
@@
We have developed a comprehensive end-to-end solution to rerank ”How-to” search query results on Android systems and applications across domains. This can be enhanced as a platform-agnostic approach with slight adaptions to support other environments, such as web or desktop. Our solution enhances search engine results by reranking pages based on a simulated verification on Android devices and preliminary experimental outcomes have demonstrated the effectiveness of our methods. Our key contributions are outlined as follows:
@@
We propose adding search result verification and reranking into the search process.
@@
For technical ”How-to” queries, we proposed a three-stage solution. At the first stage, we propose an information extraction solution to extract step-by-step instructions from web pages using generative AI with grounding techniques. At the second stage, we propose an action agent to take instructions in natural language and execute a series of actions on client device. At the third stage, we introduce a range of features that integrate both textual data and the outcomes of executions, and these features are then utilized to rerank the retrieved results.
@@
To illustrate and support future research in this direction, we developed a new research platform MagicWand 111http://nlp-in-477-l.soe.ucsc.edu:8000/admin/ that can verify and rerank search results for ”How-to” queries across several different domain applications on Android. We also collected a new dataset with human annotators.
@@ 2. Related Work
As far as we know, there is no prior work on search results verification for ”How-to” queries using simulated execution. On the other hand, the techniques we choose to use when developing different components of the proposed solution are inspired by various existing works from multiple domains. Therefore, we briefly review three important domains that influenced the technical choices we made in our research.
@@ 2.1. Information Extraction
Information extraction from web pages (Bevendorff et al., 2023), as a broader concept of our extraction task, was studied heavily back in 1990s (Finn et al., 2001). There are two methodologies: heuristics based approaches and machine learning based approaches. Heuristics approaches apply either a single metric or a set of rules along a hierarchy HTML tree to pinpoint the main content blocks. For instance, (Sun et al., 2011) considered ”text density” and the DOM tree structure, while (Pomikálek, 2011) segmented the HTML tree and then used heuristics for classification. In comparison, machine learning approaches in early studies involve sequence labeling methods and deep neural networks: (Vogels et al., 2018) employed a hidden Markov model and a convolutional neural network to classify web regions; (Jung et al., 2021) used browser-rendered visual features to enhance model performance.
Recently, researchers also leveraged LLMs to facilitate relevant downstream tasks. For example, (Gur et al., 2022) enabled T5-based models for semantic classification of HTML elements, description generation for HTML input, and autonomous web navigation. DOM-LM (Deng et al., 2022), MarkupLM (Li et al., 2021) used transformer-based architecture to represent the DOM tree and further improve classification accuracy. A new dataset PLAtE of list-like websites, such as online stores, was developed by Amazon research team to analyze how LLMs work on structure extraction (San et al., 2022). Researchers also endeavored to better utilize pre-trained LLMs to extract target information to avoid training LLMs. The instruction extraction approach used in our research is inspired by the prior work that used zero-shot GPT3 to extract plans from text (Hernandez et al., 2021).
@@ 2.2. Acting Agent
Enabling intelligent agents to behave according to instructions involves two main tasks: the first is empowering agents with sufficient knowledge to make decisions and act skillfully. The second is executing actions, collecting feedback, and recognizing environmental changes caused by these actions. To achieve the first, researchers often leverage and enhance LLMs’ reasoning and tooling capabilities (Mialon et al., 2023). For the second, agents should be enabled to interact with the host system to understand feasible actions and their results. Recent studies include using pre-trained GPT2 (Radford et al., 2019) in household simulations (Li et al., 2022a), RT1 (Brohan et al., 2022) and RT2 (Zitkovich et al., 2023) in robot control. This section will focus on how to adapt these tasks to the Android platform, on which we will carry out our experiments.
@@
An early work on Android was (Li et al., 2019) using LSTM to predict fixed action and parameter pairs on dataset RICO (Deka et al., 2017). Later, Seq2Act (Li et al., 2020) utilized Transformers for more flexible action execution given instructions. This approach set a strong baseline on a new dataset UGIF (Venkatesh et al., 2022), further enhanced by META-GUI (Sun et al., 2022) with a proposed multimodal Action Prediction Model for conversational scenarios. Along this work, Auto-UI (Zhan and Zhang, 2023) built an ”Action Chain” based model using UI representation extracted from BLIP-2 (Li et al., 2023) as visual input, which outperformed two other competitive approaches (i.e. Behavior Cloning and Simple action Chain of Thoughts (Wei et al., 2022)) on the AWIT dataset (Rawles et al., 2023). However, they overlooked the Android UI Control hierarchy, and the transformers trained without such knowledge can not predict actions accurately given ambiguous or sparse instructions. (Huang et al., 2022a) demonstrated that LLMs equipped with domain knowledge could break down high-level tasks into mid-level plans. Researchers were also starting to explore multi-modal LLM agents in Android automation, as evidenced by Yan et al. (2023); Yang et al. (2023b). Our Action Prediction Model aligns with this line of research, however, we’ve refined it by distilling the UI Control list from runtime metadata and prompting GPT4-V based on the context.
@@
Usually, researchers resorted to two technologies to detect UI Hierarchy and take actions on Android devices: One is Android Debug Bridge (ADB) (Guideline, 2023a), such as Toyama et al. (2021); Yang et al. (2023b), the other is Accessibility Service API (Guideline, 2023b), like Salehnamadi et al. (2022); Salehnamadi (2022). ADB-based solutions, like AndroidEnv (Toyama et al., 2021), adbutils (adb library for adb service, 2022), need additional connection settings for Android devices and the paired computer, which is unrealistic for end users. In comparison, Accessibility Service-based solutions can be delivered as standalone mobile applications due to the availability of Accessibility Service API on most systems. In our experiments, we used the Accessibility Service API.
@@ 2.3. Reranking
In information retrieval (IR), reranking is the adjustment of document rankings based on query-document relevance after the initial retrieval. The evolution of reranking began with RankNet (Burges et al., 2005) in 2005, applying gradient descent in the process of learning a ranking function. LambdaRank (Burges, 2010) later refined this approach by optimizing the gradients of the model scores and addressing issues caused by discontinuous gradients. And LambdaMART, combining MART (Friedman, 2001) with LambdaRank, achieved victory in the 2010 Yahoo! Learning-to-Rank Challenge. Subsequently, neural ranking models like DRMM (Guo et al., 2016) and Duet (Mitra et al., 2017) emerged. Recently, (Pobrotyn et al., 2020) introduced a context-aware neural network for item relevance calculation, and (Qin et al., 2020) proposed DASALC, a neural LTR model yielding performance comparable to LambdaMART. This progress inspires our adoption of NeuralNDCG (Pobrotyn and Bialobrzeski, 2021), which utilizes NeuralSort (Grover et al., 2019) for end-to-end, gradient-based stochastic optimization using a NDCG-based sorting operator.
@@ 3. Methods
The flowchart in Figure 1 outlines our proposed process to improve the accuracy and usefulness of ”How-to” search query results. It achieves this by extracting instructions from web pages and validating them through execution on the device. The validation outcomes are then used to refine and adjust the search result rankings. The system can overcome some limitations of traditional ranking algorithms by incorporating user-specific information (device, OS version, application version, etc.) and collecting verification results from client devices.
This is a three-stage process. Stage 1 is Instruction Extraction: given the initial top N retrieved results, a generative AI model processes each document to generate a sequence of instructions. Then, a grounding module analyzes the structure of the HTML web pages and further improves the quality of the extracted instructions. Stage 2 is On-device Execution: the Action Prediction Model reads the instructions extracted from stage 1 and predicts actions that will be taken on the device. Then, an Android agent tries to execute the actions to validate whether the instructions work with the execution information collected. This is an iterative process until the agent either completes the task or is unable to proceed further. Stage 3 is Reranking: representation features of each ¡page, query¿ pair are extracted from execution information and used for reranking the results. In this stage, search results that help yield better execution results are assigned with higher priority in the ranking.
For the simplification purpose, we have utilized pre-trained GPT models in the first two components, which enables us to rapidly establish a research platform that covers all three necessary components. Now we delve into the technical details of each component: Open Domain Instruction Extraction, On-Device Execution, and Reranking.
@@ 3.1. Open Domain Instruction Extraction
The task of extracting potentially relevant instructions from candidate web pages is divided into the generation and grounding phases, as shown in stage 1 of Figure 1. This approach enables us to extract instructions from any web page without collecting any training data, i.e. zero-shot learning.
Motivated by recent research using LLM for information extraction (Sun et al., 2011) and (Pomikálek, 2011), we first prompt a LLM to extract relevant instructions from a given web page, otherwise generate ”none”. In the generation phase, for a given search query and each candidate website , we pass in the prompt of the generative model the search query , candidate web page title, and a cleaned HTML document. The cleaned HTML is generated using a parsing algorithm that removes extraneous information, such as javascript and header information, from the web page but still preserves the hierarchical structure of the HTML snippets.
Since generative models tend to paraphrase instructions or even generate their text content unbounded with the original HTML, it is necessary to introduce a grounding mechanism to guarantee the generated instructions match the content in the HTML to minimize hallucinations in the generated content. For the grounding purpose, we first extract all text snippets , from the cleaned HTML with their corresponding XPaths. The first instruction generated from LLM is matched with the closest candidate snippet , using a combination of Faiss semantic (Johnson et al., 2019) similarity and rouge (Lin, 2004) similarity. A typical matching criterion has been illustrated in section 4.3.1. For each subsequent instruction, only snippets having similar XPaths to the previous grounded steps are considered as a matched step candidate. This approach maintains the sequential nature of extracted instructions. If a HTML page has generated instruction(s) grounded to HTML snippets, the relevant instructions can be used for the next execution stage.
@@ 3.2. On-Device Execution
The on-device execution agent is designed to directly execute instructions on an end-user device. Simulating the steps on a device provides a more accurate judgement on whether the instructions can be completed on a specific device. In our research, three components are developed to achieve this goal and the workflow is depicted in Figure 2:
Runtime UI Context Detection Module provides UI context information for the current screen state in a JSON format. It contains the instructions, UI control hierarchy, possible actions to take, and metadata for actionable regions. It serves as the input for the Action Prediction Model and is also used to extract reranking features (Sun et al., 2022).
@@
Execution Proxy Module takes the predicted action and executes it on device.

### v3

# Enhancing Mobile ”How-to” Queries with Automated Search Results Verification and Reranking

@@ Abstract.
Many people use search engines to find online guidance to solve computer or mobile device problems. Users frequently encounter challenges in identifying effective solutions from search results, often wasting time trying ineffective solutions that seem relevant yet fail to solve real problems. This paper introduces a novel approach to improving the accuracy and relevance of online technical support search results through automated search results verification and reranking. Taking ”How-to” queries specific to on-device execution as a starting point, we developed the first solution that allows an AI agent to interpret and execute step-by-step instructions in the search results in a controlled Android environment. We further integrated the agent’s findings into a reranking mechanism that orders search results based on the success indicators of the tested solutions.
@@ 1. Introduction
When lacking the knowledge to complete a task, people usually rely on a search engine to retrieve potential instructions by asking ”How-to” or similar style queries. These requests are ”procedural queries” where the user’s intent is to obtain step-by-step instructions for performing a specific task or operation. However, instructions extracted from retrieved results may not be executable, as the search engine primarily ranks pages based on the similiarty between the search query and candidate web pages (Qin and Liu, 2013). Furthermore, although user engagement information, such as search logs and page quality indicators like PageRank, are utilized for ranking results in modern search engines, the optimal solution often varies on a user-specific basis. The use of different operating systems, screen types, and app versions motivates a personalized reranker. In addition, search engines integrated with large language models (LLMs) often unavoidably involve the risk of generating inaccurate or fabricated information, commonly known as hallucinations (Lewis et al., 2020).
@@
How do users solve the problem? Typically, a user could visit multiple retrieved pages from a search engine triggered by a ”How-to” query, which briefly describes their ultimate goal. If a page that looks relevant contains some step-by-step instructions, the user might try to follow the instructions to see whether it solves the problem. However, the potential issue is that they often experience frustrations and waste considerable time attempting various instructions across different pages. Can a search engine further assist the user more efficiently? Inspired by such time-consuming manual procedure, we endeavor to substitute it with an automatic instruction verification process as an additional component of the search process. Our hypothesis is that search results verification for technical ”How-to” queries can achieve decent accuracy given the recent research progress on multimodal LLMs, especially with the help of state-of-the-art GPT.
@@
In particular, we introduce a three-stage solution to verify the instruction set and rerank the relevant pages accordingly, as illustrated in Figure 1: The first stage involves an Instruction Extraction Model to obtain step-by-step instructions from each retrieved page for a ”How-to” search query. The second stage verifies their quality by automatically simulating the instruction execution on devices. The third stage reranks retrieved results based on the execution information.
@@
As an initiative, we have developed a comprehensive end-to-end solution to rerank ”How-to” search query results on Android systems, selecting mobile applications from different domains. This can be enhanced as a platform-agnostic approach with slight adaptions to support other similar environments, such as web, iOS and desktop. Our solution enhances search engine results by reranking pages based on simulated verification on Android mobile devices, and our preliminary experimental outcomes have demonstrated the effectiveness of our methods. Our key contributions are outlined as follows:
@@
We propose adding search result verification and reranking into the search process, relieving users from tedious manual verification.
@@
For technical ”How-to” queries, we proposed a three-stage solution. In the first stage, we propose an information extraction solution to extract step-by-step instructions from web pages using generative AI with grounding techniques. In the second stage, we propose and build out a generic action agent to take instructions in natural language format and execute actions on a client device. In the third stage, we introduce a range of features based on the execution outcomes, and further utilize them to rerank the retrieved results.
@@
To support future research in this direction, we developed a new research platform MagicWand, which facilitates the verification and reranking process for ”How-to” queries of Android mobile applications. On top of it, we collected a new ”How-to” WeWeb dataset with human annotators.
@@ 2. Related Work
As far as we know, there is no prior work on search results verification for ”How-to” queries using simulated execution. Furthermore, the techniques we choose to use when developing different components of the proposed solution are motivated by existing works from multiple domains. Therefore, we briefly review three important domains that influenced the technical choices we made in our research.
@@ 2.1. Information Extraction
Information extraction from web pages (Bevendorff et al., 2023), as a broader concept of our extraction task, was studied extensively back in the 1990s (Finn et al., 2001). There are two methodologies: heuristics based approaches and machine learning based approaches. Heuristics approaches apply either a single metric or a set of rules along a hierarchy HTML tree to pinpoint the main content blocks. For instance, (Sun et al., 2011) considered ”text density” and the DOM tree structure, while (Pomikálek, 2011) segmented the HTML tree and then used heuristics for classification. In comparison, machine learning approaches in early studies involve sequence labeling methods and deep neural networks: (Vogels et al., 2018) employed a hidden Markov model and a convolutional neural network to classify web regions; (Jung et al., 2021) used browser-rendered visual features to enhance model performance.
Recently, researchers also leveraged LLMs to facilitate relevant downstream tasks. For example, (Gur et al., 2022) enabled T5-based models for semantic classification of HTML elements, description generation for HTML input, and autonomous web navigation. DOM-LM (Deng et al., 2022), MarkupLM (Li et al., 2021) used transformer-based architecture to represent the DOM tree and further improve classification accuracy. A new dataset PLAtE of list-like websites, such as online stores, was developed by Amazon research team to analyze how LLMs work in structure extraction (San et al., 2022). Researchers also endeavored to better utilize pre-trained LLMs to extract target information to avoid training LLMs. The instruction extraction approach used in our research is inspired by the prior work that used zero-shot GPT3 to extract plans from text (Hernandez et al., 2021).
@@ 2.2. Acting Agent
Enabling intelligent agents to behave according to instructions involves two main tasks: the first is empowering agents with sufficient knowledge to make decisions and act skillfully. The second is to execute actions, collect feedback, and recognize environmental changes caused by these actions. To achieve the first, researchers often leverage and enhance LLMs’ reasoning and tooling capabilities (Mialon et al., 2023). For the second, agents should be enabled to interact with the host system to understand feasible actions and their results. Recent studies include using pre-trained GPT2 (Radford et al., 2019) in household simulations (Li et al., 2022a), RT1 (Brohan et al., 2022) and RT2 (Zitkovich et al., 2023) in robot control. This section will focus on how to adapt these tasks to the Android platform on which we will carry out our experiments.
@@
An early work on Android was (Li et al., 2019) using LSTM to predict fixed action and parameter pairs on dataset RICO (Deka et al., 2017). Later, Seq2Act (Li et al., 2020) utilized Transformers for more flexible action execution given instructions. This approach set a strong baseline on a new dataset UGIF (Venkatesh et al., 2022), further enhanced by META-GUI (Sun et al., 2022) with a proposed multimodal Action Prediction Model for conversational scenarios. Along this work, Auto-UI (Zhan and Zhang, 2023) built an ”Action Chain” based model using UI representation extracted from BLIP-2 (Li et al., 2023) as visual input, which outperformed two other competitive approaches (i.e. Behavior Cloning and Simple action Chain of Thoughts (Wei et al., 2022)) on the AWIT dataset (Rawles et al., 2023). However, they overlooked the Android UI Control hierarchy, and transformers trained without such knowledge cannot accurately predict actions given ambiguous or sparse instructions. (Huang et al., 2022a) demonstrated that LLMs equipped with domain knowledge could break down high-level tasks into mid-level plans. Researchers were also starting to explore multimodal LLM agents in Android automation, as evidenced by Yan et al. (2023); Yang et al. (2023b). Our Action Prediction Model aligns with this line of research, however, we’ve refined them in a more fine-grained manner by distilling the UI Control list from runtime metadata and prompting GPT4-V dynamically based on the context.
@@
Usually, researchers resorted to two technologies to detect UI Hierarchy and take actions on Android devices: One is Android Debug Bridge (ADB) (Guideline, 2023a), such as Toyama et al. (2021); Yang et al. (2023b), the other is Accessibility Service API (Guideline, 2023b), like Salehnamadi et al. (2022); Salehnamadi (2022). ADB-based solutions, like AndroidEnv (Toyama et al., 2021), adbutils (adb library for adb service, 2022), need additional connection settings for Android devices and the paired computer, which is unrealistic for end users during daily usage. In comparison, Accessibility Service-based solutions can be delivered as standalone mobile applications due to the general availability in most systems. Considering this benefit, we decided to utilize the Accessibility Service API to develop our execution proxy, which laid down a solid foundation for action execution.
@@ 2.3. Reranking
In information retrieval (IR), reranking is the adjustment of document rankings based on query-document relevance after the initial retrieval. An important evolution of reranking began with RankNet (Burges et al., 2005) in 2005, applying gradient descent in the process of learning a ranking function. LambdaRank (Burges, 2010) later refined this approach by optimizing the gradients of the model scores and addressing issues caused by discontinuous gradients. And LambdaMART, combining MART (Friedman, 2001) with LambdaRank, achieved victory in the 2010 Yahoo! Learning-to-Rank Challenge. Subsequently, similar neural ranking models like DRMM (Guo et al., 2016) and Duet (Mitra et al., 2017) emerged. Recently, (Pobrotyn et al., 2020) introduced a context-aware neural network for item relevance calculation, and (Qin et al., 2020) proposed DASALC, a neural LTR model yielding performance comparable to LambdaMART. This progress inspires our adoption of NeuralNDCG (Pobrotyn and Bialobrzeski, 2021), which utilizes NeuralSort (Grover et al., 2019) for end-to-end, gradient-based stochastic optimization using a NDCG-based sorting operator.
@@ 3. Methods
The flow chart in Figure 1 outlines the overall process of our proposal primarily by extracting instructions from Web pages and validating them through the execution on the device. The validation outcomes are then used to refine the search result rankings. The system can overcome some limitations of traditional ranking algorithms by incorporating user-specific information (device, OS version, application version, etc.) and collecting client-side verification results.
This is a three-stage process. Stage 1 is Instruction Extraction: we take a high-level goal for an Android application as a ”How-to” search query to retrieve results from a search engine. Given the initial top N retrieved web pages, a generative AI model processes each document to generate a sequence of instructions. Then, a grounding module analyzes the structure of the HTML web pages and further improves the quality of the extracted instructions. Stage 2 is On-device Execution: the Action Prediction Model reads the instructions extracted from stage 1 and predicts actions to be taken on the device. Next, an Android agent takes those actions to validate whether the instructions work, collecting the execution information simultaneously. This is an iterative process until the agent either completes the task or can’t proceed further. Stage 3 is Reranking: representation features of each ¡page, query¿ pair are extracted from execution information and used for reranking pages for each query. In this stage, search results that help yield better execution results are assigned with higher priority during the ranking.
For simplification, we have utilized pre-trained GPT models in our implementations, which enables us to rapidly set up a pipeline that covers all three necessary components. Now we delve into the technical details of each component: Open Domain Instruction Extraction, On-Device Execution, and Reranking.
@@ 3.1. Open Domain Instruction Extraction
This stage has been divided into the generation and grounding phases, as shown in stage 1 of figure 1. Our approach allows us to extract instructions from any web page without collecting any training data, i.e. zero-shot learning 111By realizing the non-trivial effects caused by the hallucination issue, we have further refined the aforementioned strategy, making the side effect to the minimal level. More details will be covered in the discussion section, as well as our future work.
Following Sun et al. (2011); Pomikálek (2011) using LLM for information extraction, we first prompt a LLM model to extract relevant instructions from a given web page, otherwise generate ”none”. In the generation phase, for a given search query and each candidate website , we pass in the prompt of the generative model the search query , candidate web page title, and a cleaned HTML document. The cleaned HTML is generated using a parsing algorithm that removes extraneous information, such as javascript and header information, from the web page but still preserves the hierarchical structure of the HTML snippets.
Since generative models tend to paraphrase instructions or even generate text content unbounded with the original HTML, we also introduce a grounding mechanism to guarantee the generated instructions match the content in the HTML to minimize hallucinations in the generated content. For grounding, we first extract all text snippets , from the cleaned HTML with their corresponding XPaths. The first instruction generated from LLM is matched with the closest candidate snippet , using a combination of Faiss semantic (Johnson et al., 2019) similarity and rouge (Lin, 2004) similarity. A typical matching criterion has been illustrated in section 4.3.1. For each subsequent instruction, only snippets having similar XPaths to the previous grounded steps are considered as matched step candidates. This approach maintains the sequential order of extracted instructions from the original web page. If generated instruction(s) can be grounded to HTML snippets from the original web page, the relevant instructions can be used for the next execution stage.
@@ 3.2. On-Device Execution
The on-device execution agent is designed to directly execute instructions on an end-user device. Simulating the steps on a device provides more accurate judgements on whether the instructions can be completed using a specific device. In our research, three components are developed to achieve this goal with the workflow depicted in figure 2:
Runtime UI Context Detection Module delivers UI context information for the current screen state in a JSON format. It contains instructions, the UI control hierarchy, potential actions, and necessary runtime metadata. This information serves as input for the Action Prediction Model and is also used to extract reranking features (Sun et al., 2022).
@@
Execution Proxy Module takes the predicted action and executes it on the client device.